Assume z is a normal random variable with mean 0 and standard deviation 1.

1
Expert's answer
2021-12-07T13:44:15-0500

"Z\sim N(0,1)"

a)


"P(-1.55<Z<1.86)"

"=P(Z<1.86)-P(Z\leq -1.55)"

"\approx0.9685572-0.0605708"

"\approx0.907986"

b)


"P(X<-1.55\ or\ x>1.86)"

"=1-P(-1.55<X<1.86)"

"\approx1-0.907986"

"\approx0.092014"

c)


"P(Z>z)=0.015"

"1-P(Z\leq z)=0.015"

"P(Z\leq z)=0.985"

"z\approx2.1701"